OpenStudy (mattyboyy):

Jimmy's cedar trunk is in the shape of a rectangular prism and has a volume of 3/4 cubic yard. How many cubic blocks with a side length of 1/2 yard would Jimmy need to measure the volume of his trunk? A) 2 B) 3 C) 4 D) 5 E) 6

OpenStudy (igreen):

Find the volume of 1 cubic block and divide it from 3/4. \(V_{cube} = a^3\) Where 'a' is the side length. \(V_{cube} = \dfrac{1}{2}^3\)

OpenStudy (igreen):

At least, I think that's right.. :l

OpenStudy (igreen):

\(\dfrac{1}{2} \times \dfrac{1}{2} \times \dfrac{1}{2} \rightarrow \dfrac{1}{2 \times 2 \times 2}\)

OpenStudy (igreen):

2 * 2 * 2 = ?

OpenStudy (igreen):

Can you multiply 2 * 2 * 2? @mattyboyy

OpenStudy (tkhunny):

It's not that easy. The measurements of the chest are important and are not given. Example: Chest 1 yd x 1 yd x 3/4 yd